OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

-2/5

OpenStudy (saifoo.khan):

-2/5

OpenStudy (anonymous):

-2/5

OpenStudy (anonymous):

y=mx+b solve equation for y 5y = -2x - 10 > y = -2/5 x - 2 > m is slope, thus -2/5